2v+3w=60 0.5v+4w=40 i need help solving it
Answer by mukhopadhyay(490)

2v+3w=60 .......(1)
0.5v+4w=40 .....(2)
Multiply (2) with -4:
-2v -16w = -160 ....(3)
Add (3) to (1):
-13w = -100
=> w = 100/13;
From (2):
.5v+4w=40 =>.5v=40-4w => v=80-8w => v = 80-8(100/13) => v = 240/13;
Answer: (v,w) = (240/13, 100/13)